### v1

@@ 3 Basics of Gravitational Lensing: in Microlensing and Strong Lensing Regime
In the presence of matter, the propagation of GW depends on the gravitational potential of object(s) along its trajectory. The equation describing the spatial variation of the GW amplitude is given by [13],
$\left(\nabla^{2}+\frac{\omega^{2}}{c^{2}}\right)\tilde{\phi}(\vec{r},\omega)=\frac{4\omega^{2}U(\vec{r})}{c^{4}}\tilde{\phi}(\vec{r},\omega),$ (3.1)
@@
where $\nabla^{2}$ is the 3-dimensional Laplacian operator, $\omega$ is the angular frequency ($=2\pi f$) of the gravitational wave, $U(\vec{r})$ is the gravitational potential (assumed not to be varying in the scale of the passing time of the GW) of the lens and the GW strain in the frequency domain is given by, $h_{\mu\nu}(\omega,\vec{r})=e_{\mu\nu}\tilde{\phi}(\omega,\vec{r})$ where $\tilde{\phi}(\omega,\vec{r})$ consists of amplitude and phase of the GW and $e_{\mu\nu}$ is the associated polarization.
@@
Figure 2: The figure shows a schematic diagram for the gravitational lensing with a thin lensing approximation. The deflection of the GW happens only at the lens plane.
@@
The thin lens approximation assumes that the deflection of the GW caused by the lens happens at the lens plane, a plane passing through the center of the lens at which the lensing object in projected. In figure 2, we have showed a ray diagram schematic for gravitational lensing. Let, $D_{l}$ be the angular diameter distance  [^2] Angular diameter distance is a distances inferred from the angular size of the sources. See [58] between the observer and the lens, $D_{s}$ is the angular diameter distance between the observer and the source and $D_{ls}$ is the angular diameter distance between the lens and the source. $\vec{\eta}$ is the position of the source in the source plane. The source plane origin is located at the intersection the extended straight line joining the the observer to the center of the lens. $\vec{\xi}$ is the impact parameter of the incoming ray in the lens plane. The origin of the lens plane is located at the center of the lens. We have defined two dimensionless vectors $\vec{x}$ and $\vec{y}$ as following:
$\vec{x}=\frac{\vec{\xi}}{\xi_{0}}\quad\text{and}\quad\vec{y}=\frac{D_{l}\vec{\eta}}{\xi_{0}D_{s}},$ (3.2)
where $\xi_{0}$ sets the characteristic distance scale of the system, called the Einstein radius.
@@
The thin lens approximation is valid when the dimension of the lens is much smaller than the total travelled distance by the GW. The strength of amplification leads to the classification of gravitational lensing as strong lensing or micro-lensing. The amplification factor is defined as $A(f)=\frac{\tilde{\phi}^{\rm lensed}(f,\vec{r})}{\tilde{\phi}^{\rm unlensed}(f,\vec{r})}$. The term $\tilde{\phi}^{\rm lensed}(f,\vec{r})$ at the numerator, is dependent on the relative time delay between the arrival of GWs through different spatial trajectories. This time delay includes the geometric time delay between GWs travelling from different deflected trajectories and the Shapiro time delay, caused by slowing of time at higher spacetime curvature regions. The total time delay is given by [13, 50],
$t_{d}(\vec{x},\vec{y})=\dfrac{D_{s}\xi_{0}^{2}}{cD_{l}D_{ls}}\left(1+z_{l}\right)\left[\frac{1}{2}|\vec{x}-\vec{y}|^{2}-\psi(\vec{x})+\Phi_{m}(\vec{y})\right]\quad,$ (3.3)
@@
where $z_{l}$ is the redshift of the lens, obtained from its angular diameter distance, $c$ be the speed of light in vacuum. The first term provides the geometric time-delay and the second term provides the Shapiro time-delay where $\psi(\vec{x})$ provides the gravitational potential of the lens in the lens plane. The third term with $\Phi_{m}(\vec{y})$ is optional, to set the minimum $t_{d}$ value to zero. In presence of gravitational lensing, GW waves are time-delayed and deflected therefore the addition of GWs in different phases causes interference of lensed GW in the wave-optics limit.
@@
Figure 3: The figure shows the modulus of the amplification factor for BH lens and SIS lens for three different values of y: y=2, y=3 and y=5, in the range $w\in[10^{-2},5]$
@@
$A(f)=\frac{D_{s}\xi_{0}^{2}}{cD_{l}D_{ls}}\frac{f}{i}\int d^{2}\vec{x}\exp\left[2\pi ift_{d}(\vec{x},\vec{y})\right].$ (3.4)
@@
here $J_{0}$ is the spherical Bessel function of order zero, and $w$ is the dimensionless frequency defined as, $w=\dfrac{8\pi GM_{lz}f}{c^{3}}$ where $M_{lz}=M_{l}(1+z_{l})$ is the redshifted mass of the lens, $f$ be the frequency of the GW and G is the gravitational constant.
Different lensing regimes can also be described by the dimensionless parameter $w$. When $w\leq 1$, we are in the wave optics (microlensing) regime, the amplification factor amplitude $|A(w)|$ and its phase $\theta_{A}(w)=-i\rm{log_{e}}\left(\frac{A(w)}{\lvert A(w)\rvert}\right)$ are very oscillatory in nature. However, in the limit $w\gg 1$, we are in the geometric optics (strong lensing) regime. Both $|A(w)|$ and $\theta_{A}(w)$ converge to a single value independent of the frequency of the gravitational waves. The amplification factor in the strong lensing regime (for any generic mass or size of lens) becomes,
$F(f)=\sum_{j}\left|\mu_{j}\right|^{1/2}\exp\left[2\pi ift_{d,j}-i\pi n_{j}\right],$ (3.6)
where, the flux-magnification of the j-th image is $\mu_{j}=1/\operatorname{det}\left(\frac{\partial\vec{y}}{\partial\vec{x}_{j}}\right)$, and $t_{d,j}=t_{d}(\vec{x_{j}},\vec{y})$ and $n=0,1/2,1$ for minimum, saddle point, maximum of the $t_{d}(\vec{x},\vec{y})$ function. They are known as type-I, type-II and type-III images respectively. Together with these. we can write the lensed wave amplitude in the time-domain is then represented as
@@
Figure 4: The figure shows the capability of our model in capturing the necessary features present in the microlensing amplification.
$\phi^{\rm lensed}(t,\vec{r})=\sum_{j}\left|\mu_{j}\right|^{1/2}\phi^{\rm unlensed}\left(t-t_{d,j},\vec{r}\right)\exp\left[-i\pi n_{j}\right]$ (3.7)
@@
For the black hole lens and singular isothermal sphere profile lens, we have analytic solutions [59]. We plot the solutions as a function of w with y as a parameter in figure 3. However, such analystical solution may not exist for more general lens including assymetric shape and baryonic effects. As a result, for lensing search pipeline to discover signal, it is important for it to capture any general lensing scenario. In a specific lensing model agnostic approach, we have constructed frequency modulated template that captures most (if not all) of the features of microlensing of gravitational waves. The model is described as:
$A(f)=a\left[1+be^{-kf}\cos\left(\frac{2\pi f}{f_{0}}+\phi\right)\right],$ (3.8)
@@
where $a$ is microlensing amplitude  [^3] If strong lensing present together with microlensing, it would also contribute to this amplitude, $a=a_{strong}\times a_{micro}\equiv\sqrt{\mu}\times a_{micro}$, $b$ captures perturbation scale at which the microlensing in the amplification varies, $f_{0}$ denotes characteristic frequency of the oscillations in amplification, $\phi$ denotes the phase shift due to lensing, and $e^{-kf}$ ensures that the amplification slowly converges to $a$ as $f$ is much greater than $f_{0}$. In the later sections, we use this amplification model to understand the different aspects of a microlensing detection and characterization. In figure 4, we have demonstrated the capability to contain all the features in a microlensing amplification for the case with $y=1$ with a BH lens and an SIS lens (With lens of mass $500\rm M_{\odot}$ at a redshift of $z_{l}=0.2$) . The fitting parameters for the black hole amplification profile are $(a,b,k,f_{0},\phi)=(1.113,0.322,-0.0008,40.383\,\rm{Hz},4.613)$ and the fitting parameters for the SIS amplification profile are $(a,b,k,f_{0},\phi)=(1.421,0.206,0.002,41.565\,\rm{Hz},4.745)$  [^4] Considering no strong lensing $\sqrt{\mu}=1$, these fitting parameters have been obtained.. We observe that, the amplification model with appropriate fitting parameters can nicely mimic the microlensing amplification [^5] In the current setup, the microlensing model does not capture the phase modulation. But it can be trivially included by a few additional variables to capture the phase modulations, in expense of computational additional time. Thus to demonstrate this technique, we have taken a model with amplitude modulation..
@@ 4 Methodology
$R_{i}(t)\equiv d_{i}(t)-h_{BFi}(t)=r_{i}(t)+n_{i}(t),$ (4.1)
here $r_{i}(t)$ denotes the residual signal and $n_{i}(t)$ denotes the noise present in the data. We define the cross-correlation between two residual data as from two detectors denoted by $\{x,x^{\prime}\}$,
[…]
$=S_{xx^{\prime}}(t)+N_{xx^{\prime}}(t)+P_{xx^{\prime}}(t)+Q_{xx^{\prime}}(t),$
$\approx S_{xx^{\prime}}(t)+N_{xx^{\prime}}(t),$ (4.2)
where $\tau$ is the timescale over which the cross-correlation is performed and $t_{d}$ be the time-delay between the arrival of the signal (thus the residual) between the detectors. The second expression contains four terms are : $S_{xx^{\prime}}\equiv r_{x}\otimes r_{x^{\prime}}$, $N_{xx^{\prime}}\equiv n_{x}\otimes n_{x^{\prime}}$, $P_{xx^{\prime}}\equiv r_{x}\otimes n_{x^{\prime}}$, $Q_{xx^{\prime}}\equiv n_{x}\otimes r_{x^{\prime}}$ respectively, where $\otimes$’ denotes the cross-correlation between them. The final expression is approximated when the cross-correlation timescale ($\tau$) is sufficiently large, comparable to the duration of the signal, the cross-terms do not contribute much.
Therefore, if the absence of any microlensing, the best-fit would be able to capture the GW waveform, making the cross-correlation signal $S_{xx^{\prime}}$ negligible. However, if microlensing present in the GW signal, all its features are not captures by the best-fit signal, therefore it shows up in the cross-correlation. In this analysis, we have ignored waveform systematics and its impact on lensing. The cross-correlation technique can find any non-general relativistic unmodelled signatures present in the signal as well [60].
A similar but distinct detection approach has been taken with strong lensing regime when there are multiple images of a GW signal separated by time delays from seconds to years. We have showed in our previous work GLANCE [52] that, such strongly lensed GW signals can be detected by performing cross-correlation between the reconstructed one-polarization signals of the images using a detector pair at two different times. We defined lensing SNR to quantify how much the cross-correlation signal differs from the noise cross-correlation values. The method demostrates that for type-I and type-II pair (or type-II and type-III pair) of images, the cross-correlation shows up by among the different polarizations, otherwise it shows up in the same polarization. A joint parameter estimation was performed to explore the degeneracies between the source and lens parameters and thus helping in removing the lensing bias. Therefore to look for the strong lensing aspects from the data, we have polarization cross-correlation based technique GLANCE and to look at the microlensing aspects, we have now developed residual cross-correlation based technique $\mu$-GLANCE.
@@ 5 Application of $\mu$-GLANCE on simulated GW data
As a demonstration of the technique, we generated GW data using a system of binary black holes with masses $35M_{\odot}$ and $25M_{\odot}$ at $1$ Gpc luminosity distance with inclination angle $0.5$ and dimensionless aligned spins of the primary and the secondary as $0.1$ and $0.2$ respectively. We generate the waveform using the phenomenological model IMRPhenomXPHM [55] and used the analytic noise power spectral density (PSD) model AdVO4T1800545 and AdvVirgo  [^6] See [61] and [62] respectively to generate the noise for the LIGO and Virgo detectors respectively. We choose the amplification factor  [^7] A more generic approach consists of a phase $\phi$ in the cosine term of the amplification factor and estimating that parameter as well, but this form of the amplification bears all the necessary modifications required. In the appendix A, we have showed that the exclusion of the phase $\phi$ does not change any significant results performed in this analysis. as
$A(f;a,b,k,f_{0})=a\left[1+be^{-kf}cos\left(2\pi\frac{f}{f_{0}}\right)\right].$ (5.1)
@@
The value of the strong lensing parameter $\sqrt{\mu}\equiv a$ is chosen to be $1.8$ with microlensing perturbation parameter $b=0.4$, microlensing oscillation scale parameter $f_{0}=30$ Hz and the microlensing convergence parameter $k=0.004$. We obtained the best-fit strain without the hypothesis of microlensing present in the data. The data, injected signal and the best-fit (obtained in the frequency domain and then reverted back in time domain) is shown in the figure 5. The contribution of the GW observatory KAGRA is not discussed in this work, since the KAGRA sensitivity is less in comparison to the currently ongoing fourth observation run sensitivity of LIGO and Virgo detectors and therefore inclusion of KAGRA will not make any significant differences in the results.
@@
Figure 5: The figure shows the data, signal injected and the best-fit signal for the LIGO-H, LIGO-L and Virgo-V observatories. The best-fit is obtained first by performing an estimation of the source parameters (masses, aligned spins, luminosity distance and inclination angle) and then using the medians of the posterior distributions of each.
@@
We calculate the residual at each detector by subtracting the best-fit signal from the data. We align the residuals in time by considering the RA and Dec distributions of the GW event in the skymap. We smoothen the data by passing it through a bandpass with lower cut-off and upper cut-off frequencies respectively at $30$ Hz and $512$ Hz  [^8] The choice of the band is very specific to this example, with lower mass systems, we may need to set up the upper cut-off at higher frequencies. In figure 6, we have shown the residual cross-correlation in the presence of the GW microlensing signal and in the absence of the signal, for pairs formed within LIGO-H, LIGO-L and Virgo-V. The signal is injected close to the fourth second  [^9] The merger happens at the 3.9th second of the signal. in a $8$ second long time-series data, that is why we observe a cross-correlation peak at that position. The cross-correlation timescale $\tau=0.5s$, since that is the typical duration of the signal. It is noteworthy to mention that, the match-filtering signal to noise ratio (SNR) for the residual drops around 70-80% when compared with the match-filtering SNR obtained from the data.
@@
Figure 6: The figure shows the cross-correlation signal between microlensed GW residuals for three detector configuration using H, L, and V, when the GW signal is present ($D_{xx^{\prime}}$) and when the GW signal is not present ($N_{xx^{\prime}}$). The cross-correlation timescale is chosen as $\tau=0.5s$.
@@
Figure 7: Residual cross-correlation with H, L and V residual strains as opposed to the noise cross-correlation distribution. The cross-correlation timescale chosen here is $\tau=0.5s$. The choice of $\tau$ is decided by the total mass of the system, with lower $\tau$’s chosen for higher total mass systems. The lines from
@@
Figure 8: The figure shows the variation of the residual cross-correlation $D_{xx^{\prime}}$ with cross-correlation timescale. We observe that as the cross-correlation timescale is increased, the noise-fluctuations go down, making the peak of the residual cross-correlation prominent.

### v2

@@ 3 Basics of Gravitational Lensing: in Wave-optics and Geometric-optics lensing Regimes
In the presence of matter, the propagation of GW depends on the gravitational potential of object(s) along its trajectory. The equation describing the spatial variation of the GW amplitude is given by (Schneider et al., 1992),
$\left(\nabla^{2}+\frac{\omega^{2}}{c^{2}}\right)\tilde{\phi}(\vec{r},\omega)=\frac{4\omega^{2}U(\vec{r})}{c^{4}}\tilde{\phi}(\vec{r},\omega),$ (1)
@@
where $\nabla^{2}$ is the 3-dimensional Laplacian operator, $\omega$ is the angular frequency ($=2\pi f$) of the GW, $U(\vec{r})$ is the gravitational potential (assumed not to be varying in the scale of the passing time of the GW) of the lens and the GW strain in the frequency domain is given by, $h_{\mu\nu}(\omega,\vec{r})=e_{\mu\nu}\tilde{\phi}(\omega,\vec{r})$ where $\tilde{\phi}(\omega,\vec{r})$ consists of amplitude and phase of the GW and $e_{\mu\nu}$ is the associated polarization.
@@
Figure 2: The figure shows a schematic diagram for the gravitational lensing with a thin lensing approximation. The thin lens approximation assumes the deflection of the ray happens only at the lens plane.
@@
The thin lens approximation assumes that the deflection of the GW caused by the lens happens at the lens plane, a plane passing through the centre of the lens at which the lensing object in projected. In figure 2, we have showed a ray diagram schematic for gravitational lensing. Let, $D_{l}$ be the angular diameter distance  [^2] Angular diameter distance is a distances inferred from the angular size of the sources. See (Hogg, 1999) between the observer and the lens, $D_{s}$ is the angular diameter distance between the observer and the source and $D_{ls}$ is the angular diameter distance between the lens and the source. $\vec{\eta}$ is the position of the source in the source plane. The source plane origin is located at the intersection the extended straight line joining the the observer to the centre of the lens. $\vec{\xi}$ is the impact parameter of the incoming ray in the lens plane. The origin of the lens plane is located at the centre of the lens. We have defined two dimensionless vectors $\vec{x}$ and $\vec{y}$ as following:
where $\xi_{0}$ sets the characteristic distance scale of the lens system.
@@
The thin lens approximation is valid when the dimension of the lens is much smaller than the total travelled distance by the GW. The depth of gravitational potential which in turn controls the strength of the amplifications leads to the classification of gravitational lensing as strong lensing or microlensing. The amplification factor is defined as $A(f)=\frac{\tilde{\phi}^{\rm lensed}(f,\vec{r})}{\tilde{\phi}^{\rm unlensed}(f,\vec{r})}$. The term $\tilde{\phi}^{\rm lensed}(f,\vec{r})$ at the numerator, is dependent on the relative time delay between the arrival of GWs through different spatial trajectories. This time delay includes the geometric time delay between GWs travelling from different deflected trajectories and the Shapiro time delay, caused by slowing of time at higher spacetime curvature regions. The total time delay is given by (Schneider et al., 1992; Takahashi & Nakamura, 2003),
[…]
@@
where $z_{l}$ is the redshift of the lens corresponding the angular diameter distance of the lens $D_{l}$ , $c$ be the speed of light in vacuum. The first term provides the geometric time-delay and the second term provides the Shapiro time-delay where $\psi(\vec{x})$ provides the gravitational potential of the lens in the lens plane. The third term with $\Phi_{m}(\vec{y})$ is optional, to set the minimum $t_{d}$ value to zero. In presence of gravitational lensing, GW waves are time-delayed and deflected therefore the addition of GWs in different phases causes interference of lensed GW in the wave-optics limit.
@@
Figure 3: The figure shows the modulus of the amplification factor for point-mass lens and SIS lens for three different values of y: y=0.5, y=1.0 and y=2.0, in the range $w\in[0.1,10]$
@@
$A(f)=\frac{D_{s}\xi_{0}^{2}}{cD_{l}D_{ls}}\frac{f}{i}\int d^{2}\vec{x}\exp\left[2\pi ift_{d}(\vec{x},\vec{y})\right].$ (4)
@@
[…]
$A(f)=\sum_{j}\left|\mu_{j}\right|^{1/2}\exp\left[2\pi ift_{d,j}-i\pi n_{j}\right],$ (6)
[…]
@@
Figure 4: The figure shows the capability of the amplification model in capturing the necessary features present in the WO microlensing amplification. We picked up two cases from the figure 3 and fitted it with our generic amplification model. The fitting parameters are mentioned in the text.
$\phi^{\rm lensed}(t,\vec{r})=\sum_{j}\left|\mu_{j}\right|^{1/2}\phi^{\rm unlensed}\left(t-t_{d,j},\vec{r}\right)\exp\left[-i\pi n_{j}\right]$ (7)
@@
For the point mass lens and SIS lens, we have analytic solutions (Tambalo et al., 2023). We plot the solutions as a function of $w$ with $y$ as a parameter in figure 3. However, such analytical solution may not exist for more general lens including asymmetric shape and non-spherical mass distribution. Also, the impact of baryons on the lens profile, and consequently on the WO microlensing signature, remains largely uncertain. As a result, for lensing search pipeline to discover signal with wave-optics effects, it is important for it to capture any general lensing scenario. We note that any wave-optics effect is captured by a scale close to unity and constant periodicity in the frequency-domain (both captured by the misalignment parameter $y$ and the dimensionless frequency $w\propto M_{lz}f$). Also these wave effects decreases as $w$ becomes much larger than unity, this is the geometric-optics regime. Driven by these observations we follow a lens-model agnostic approach, and construct a frequency modulated template that captures most (if not all) of the features of WO microlensing of GWs. The model is described as:
$A(f)=a\left[1+be^{-kf}\cos\left(\frac{2\pi f}{f_{0}}+\phi\right)\right],$ (8)
@@
where $a$ is lensing amplitude  [^3] If geometric-optics (GO) lensing present together with WO microlensing, it would also contribute to this amplitude, $a=a_{GO}\times a_{WO}\equiv\sqrt{\mu}\times a_{WO}$, $b$ captures perturbation scale at which the WO microlensing in the amplification varies, $f_{0}$ denotes characteristic frequency of the oscillations in amplification, $\phi$ denotes the phase shift due to lensing, and $e^{-kf}$ ensures that the amplification slowly converges to $a$ as $f$ is much greater than $f_{0}$. In the later sections, we use this amplification model to understand the different aspects of a WO microlensing detection and characterization. In figure 4, we have demonstrated the capability to contain all the features in a WO microlensing amplification for the case with $y=1$ with a point-mass lens and an SIS lens (With lens of mass $500\rm M_{\odot}$ at a redshift of $z_{l}=0.2$) . The fitting parameters for the point-mass amplification profile are $(a,b,k,f_{0},\phi)=(1.113,0.322,-0.0008,40.383\,\rm{Hz},4.613)$ and the fitting parameters for the SIS amplification profile are $(a,b,k,f_{0},\phi)=(1.421,0.206,0.002,41.565\,\rm{Hz},4.745)$  [^4] Considering no geometric-optics lensing $\sqrt{\mu}=1$, these fitting parameters have been obtained.. We observe that, the amplification model with appropriate fitting parameters can nicely mimic the WO microlensing amplification [^5] In the current setup, the WO microlensing model does not capture the phase modulation. But it can be trivially included by a few additional variables to capture the phase modulations, in expense of computational additional time. Thus to demonstrate this technique, we have taken a model with amplitude modulation..
@@ 4 Methodology
$R_{i}(t)\equiv d_{i}(t)-h_{BFi}(t)=r_{i}(t)+n_{i}(t),$ (9)
here $r_{i}(t)$ denotes the residual signal and $n_{i}(t)$ denotes the noise present in the data. We define the cross-correlation between two residual data as from two detectors denoted by $\{x,x^{\prime}\}$, {align} D_x x’ (t) &≡R_x ⊗R_x’ = 1τ ∫_t-τ/2 ^t+τ/2 r_x (t’) r_x’ (t’ + Δt_det) dt’, = S_x x’(t) + N_x x’(t) + P_x x’(t) + Q_x x’(t), ≈S_x x’(t) + N_x x’(t), where $\tau$ is the timescale over which the cross-correlation is performed and $\Delta t_{\rm det}$ be the time-delay between the arrival of the signal (thus the residual) between the detectors. The second expression contains four terms are : $S_{xx^{\prime}}\equiv r_{x}\otimes r_{x^{\prime}}$, $N_{xx^{\prime}}\equiv n_{x}\otimes n_{x^{\prime}}$, $P_{xx^{\prime}}\equiv r_{x}\otimes n_{x^{\prime}}$, $Q_{xx^{\prime}}\equiv n_{x}\otimes r_{x^{\prime}}$ respectively, where $\otimes$’ denotes the cross-correlation between them. The final expression is approximated when the cross-correlation timescale ($\tau$) is sufficiently large, comparable to the duration of the signal, the cross-terms do not contribute much.
Therefore, if the absence of any WO microlensing, the best-fit would be able to capture the GW waveform, making the cross-correlation signal $S_{xx^{\prime}}$ negligible. However, if WO microlensing present in the GW signal, all its features are not captures by the best-fit signal, therefore it shows up in the cross-correlation. In this analysis, we have ignored waveform systematics and its impact on lensing. The cross-correlation technique can find any non-general relativistic unmodelled signatures present in the signal as well (Dideron et al., 2023).
A similar but distinct detection approach has been taken with strong lensing regime when there are multiple images of a GW signal separated by time delays from seconds to years. We have showed in our previous work GLANCE (Chakraborty & Mukherjee, 2024) that, such strongly lensed GW signals in the geometric-optics limit can be detected by performing cross-correlation between the reconstructed one-polarization signals of the images using a detector pair at two different times. We defined lensing SNR to quantify how much the cross-correlation signal differs from the noise cross-correlation values. The method demonstrates that for type-I and type-II pair (or type-II and type-III pair) of images, the cross-correlation shows up by among the different polarizations, otherwise it shows up in the same polarization. A joint parameter estimation was performed to explore the degeneracies between the source properties and the lensing effects and thus helping in removing the lensing bias. Therefore to look for the geometric-optics lensing aspects from the data, we have polarization cross-correlation based technique GLANCE and to look at the WO microlensing aspects, we have now developed residual cross-correlation based technique $\mu$-GLANCE.
@@ 5 Application of $\mu$-GLANCE on simulated GW data
As a demonstration of the technique, we generated GW data using a system of binary black holes with masses $35M_{\odot}$ and $25M_{\odot}$ at $1$ Gpc 6661 Gpc = $10^{9}$ parsec (pc), where 1 pc = 3.086e+16 metre luminosity distance with inclination angle $0.5$ and dimensionless aligned spins of the primary and the secondary as $0.1$ and $0.2$ respectively. We generate the waveform using the phenomenological model IMRPhenomXPHM (Pratten et al., 2021) and used the analytic noise power spectral density (PSD) model AdVO4T1800545 and AdvVirgo  [^7] See (Barsotti et al., 2018) and (Manzotti & Dietz, 2012) respectively to generate the noise for the LIGO and Virgo detectors respectively. We choose the amplification factor  [^8] A more generic approach consists of a phase $\phi$ in the cosine term of the amplification factor and estimating that parameter as well, but this form of the amplification bears all the necessary modifications required. In the appendix A, we have showed that the exclusion of the phase $\phi$ does not change any significant results performed in this analysis. as
$A(f;a,b,k,f_{0})=a\left[1+be^{-kf}cos\left(2\pi\frac{f}{f_{0}}\right)\right].$ (10)
@@
The value of the strong lensing magnification $\sqrt{\mu}\equiv a$ is chosen to be $1.8$ with WO microlensing perturbation parameter $b=0.4$, WO microlensing oscillation scale parameter $f_{0}=30$ Hz and the microlensing WO to geometric-optics transition parameter $k=0.004$. We obtained the best-fit strain without the hypothesis of WO microlensing present in the data. The data, injected signal and the best-fit (obtained in the frequency domain and then reverted back in time domain) is shown in the figure 5. The contribution of the GW observatory KAGRA is not discussed in this work, since the KAGRA sensitivity is less in comparison to the currently ongoing fourth observation run sensitivity of LIGO and Virgo detectors and therefore inclusion of KAGRA will not make any significant differences in the results.
@@
Figure 5: The figure shows the data, signal injected and the best-fit signal for the LIGO-H, LIGO-L and Virgo-V observatories. The best-fit is obtained first by performing an estimation of the source parameters (masses, aligned spins, luminosity distance and inclination angle) with an unlensed hypothesis and then using the medians of the posterior distributions.
@@
We calculate the residual at each detector by subtracting the best-fit signal from the data. We align the residuals in time by considering the RA and Dec distributions of the GW event in the sky-map. We smoothen the data by passing it through a bandpass with lower cut-off and upper cut-off frequencies respectively at $30$ Hz and $512$ Hz  [^9] The choice of the band is very specific to this example, with lower mass systems, we may need to set up the upper cut-off at higher frequencies. In figure 6, we have shown the residual cross-correlation in the presence of the WO microlensed GW signal and in the absence of the signal, for pairs formed within LIGO-H, LIGO-L and Virgo-V. The signal is injected close to the fourth second  [^10] The merger happens at the 3.9th second of the signal. in a $8$ second long time-series data, that is why we observe a cross-correlation peak at that position. The cross-correlation timescale $\tau=0.5s$, since that is the typical duration of the signal. It is noteworthy to mention that, the matched-filtering signal to noise ratio (SNR) for the residual drops around 70-80% when compared with the matched-filtering SNR obtained from the data.
@@
Figure 6: The figure shows the cross-correlation signal between WO microlensed GW residuals for three detector configuration using H, L, and V, when the GW signal is present ($D_{xx^{\prime}}$) and when the GW signal is not present ($N_{xx^{\prime}}$). The cross-correlation timescale is chosen as $\tau=0.5s$.
@@
Figure 7: The figure shows the residual cross-correlation with H, L and V residual strains as opposed to the noise cross-correlation distribution. The cross-correlation timescale chosen here is $\tau=0.5s$. The choice of $\tau$ is decided by the total mass of the system, with lower $\tau$’s chosen for higher total mass systems. The lines from show the strength of the cross-correlation and the histogram is constituted from the noise cross-correlation values.
@@
Figure 8: The figure shows the variation of the residual cross-correlation $D_{xx^{\prime}}$ with cross-correlation timescale. The noise cross-correlation values follow a random distribution with a mean tending towards zero. We observe that as the cross-correlation timescale is increased, the noise-fluctuations go down, making the the residual cross-correlation more prominent.